Solve the problem.
Your company uses the quadratic model y = -11x2 + 350x to represent how many units (y) of a new product will be sold (x) weeks after its release. How many units can you expect to sell in week 17?
I would appreciate help on this.Thanks.

Found 2 solutions by RAY100, nerdybill:
Answer by RAY100(1637)

y=-11x^2 +350x
.
y(17) = -11*(17)^2 +350*17
.
y(17) =-3179 + 5950
.
y(17) = 2771
.

Answer by nerdybill(7384)

Set x=17
into
y = -11x^2 + 350x
and solve:
y = -11(17)^2 + 350(17)
y = -3179 + 5950
y = 2771